# A typical medium cluster approach for multi-branch phonon localization

### v1

@@ I Introduction
Various numerical methods including continuum field theory John et al. (1983), diagrammatic technique Kirkpatrick (1985), exact diagonalization Payton and Visscher (1967); Monthus and Garel (2010), multifractal analysis Ludlam et al. (2003), transfer matrix method Pinski et al. (2012) have been put forth to investigate the phonon localization, especially in simplified models. Alternatively, effective medium methods, such as Coherent Potential Approximation (CPA) Soven (1967); Elliott et al. (1974). and its cluster extensions, including the Dynamical Cluster Approximation (DCA)  Jarrell and Krishnamurthy (2001); Ekuma et al. (2014); Terletska et al. (2014); Ekuma et al. (2015); Zhang et al. (2015, 2016); Tam et al. (2021) have been used to model simple phonon systems. The random binary alloys with mass and force-constant disorder and with the vector nature of vibrations have been studied using the itinerant coherent-potential approximation (ICPA ) Ghosh et al. (2002). However, this method in its present formalism is not able to explore the localization aspect of the phonons. Also, in this context, various first-principles-based calculations Magdău and Ackland (2017); Zhang et al. (2022) have recently been utilized to study phonon localization in realistic systems.
@@
Despite the progress, modeling, and simulation of complex realistic features responsible for phonon localization in materials remains challenging. An effective numerical method would involve constructing material-specific force-constant models and solving them. The absence of lattice translational symmetry in disordered solids and the breakdown of the Bloch theorem make this task difficult to solve straightforwardly. Moreover, force-constants calculations are computationally expensive compared to electronic case hopping calculations, and maintaining the force-constants sum rule can be challenging, especially in the presence of disorder. The inclusion of vibration directions adds another layer of complexity to the problem due to the correlations, where the vibration of an atom at site $R_{l}$ in the $\alpha$ direction is coupled with the vibration of an atom at site $R_{l^{\prime}}$ in the $\beta$ direction. Furthermore, accurately capturing Anderson localization in such systems is challenging, as it often requires the calculation of two-particle transport quantities or finite-size scaling analysis that can be computationally expensive, particularly for complex realistic models.  Zhang et al. (2017)
@@
In this paper, to enable a more realistic modeling of phonon systems and their localization within the effective medium framework, we have developed the multi-branch cluster DCA and its typical medium variant, TMDCA. The DCA and TMDCA, are two cluster-effective medium methods that were originally developed for electronic systems Jarrell and Krishnamurthy (2001); Ekuma et al. (2014); Terletska et al. (2014); Ekuma et al. (2015); Zhang et al. (2015, 2016); Tam et al. (2021), and are used to study the disorder effects and localization, respectively. In the present work, building upon our previous work on phonon localization in a mass-disordered binary alloy using a single-direction vibration model Mondal et al. (2017), we have extended the cluster formalism to three directions of vibrations We validate the developed formalism against several exact diagonalization results and limiting scenarios. Applying our developed multi-branch formalism, we investigate how the vector nature of vibrations affects the spectra and localization of phonons. After comparing the results of single-branch and multi-branch situations, we found that the Anderson transition of phonons is not significantly influenced by the vector nature of the phonons for the specific model considered in this work. Overall, the present study provides further understanding of the localization of vector phonons and brings the TMDCA one step closer to being an efficient numerical tool for studying Anderson localization of phonons in materials from first principles.
@@ B. Green’s function formalism
[…]
@@ 4. We then construct the cluster Green’s function for a given disorder configuration $V$:
here $\mathcal{D}$, ${D}$ are $N_{b}N_{c}\times N_{b}N_{c}$ matrices in the real space of the cluster. In our analysis, we consider a binary alloy disorder described by the following distribution function
@@
where <>expectation<...>< … > indicates an averaging over disorder configurations.
@@ D. Typical medium dynamical cluster approximation for phonons
[…]
@@ 2.1 Multi-branch force constants model
$\Phi^{\alpha\beta}(\vec{R}_{l},\vec{R}_{l})$ $-6\times\tau$
@@ A. Benchmarking of the DCA and ED methods
We start the discussion of our results by first comparing our developed multi-branch cluster DCA method with the ED results for a binary isotopic alloy system in three dimensions. This benchmarking analysis aims to evaluate the accuracy and effectiveness of our developed multi-branch cluster DCA approach in capturing the system’s behavior. For this, in Fig. 1, we compare the average density of states (ADOS) calculated for an impurity concentration of $c_{A}=0.2$ and a disorder potential of $V=0.9$, for a fixed value of $\tau=0.3$ and varying values of $\varepsilon$. First, on panel Fig. 1-a), we consider the case of $\varepsilon=0.0$, where we compare the DCA and ED results, and observe a perfect agreement between two methods. For the finite values of $\tau$, we anticipate the emergence of two impurity modes in the higher frequency regions. Specifically, we expect a doubly degenerate impurity band at the lower end of the high-frequency region and a single impurity band at the higher end, corresponding to the eigenvalues of the $3\times 3$ onsite matrix $\Phi^{\alpha\beta}(l,l)$: $1-\tau,1-\tau,1+2\tau$ as explained in section  II. As shown in Fig. 1, the two impurity modes around frequency regions $42<\omega<$ 4 and $64<\omega<$ 6 are well captured by our DCA method and are in perfect agreement with the ED results.
@@ C. $\varepsilon\neq 0$ and $\tau\neq 0$: inter-branch coupling case
In Fig. 3, to demonstrate the impact of inter-branch coupling on phonon localization, we plot the ADOS and TDOS for increasing values of $\varepsilon=0.0,0.2,0.4$ at a fixed disorder potential $V_{A}=0.9$ and small impurity concentration of $c_{A}=0.2$. Beginning with the branch-decoupled case ($\varepsilon=0.0$), as shown in Fig. 3-a), we observe strong localization for high-frequency phonon modes within the frequency range of $\omega=2.5$ to $\omega=6.0$. This is evident from the TDOS approaching zero while the ADOS remains finite. The results for finite $\varepsilon=0.2$ and $\varepsilon=0.4$ are shown in Fig. 3-b) and Fig. 3-c). The data indicate a negligible effect of $\varepsilon$ on the localized phonons, although subtle changes are observed in the host modes. Specifically, the host mode becomes slightly broader and more delocalized with increasing $\varepsilon$. However, high-frequency phonons remain localized even as $\varepsilon$ increases. This is clearly illustrated in Fig. 3-d), where the ratio of TDOS/ADOS remains nearly constant across the range of considered $\varepsilon$ values (00 to $0.4$).
@@
Next, we consider an opposite scenario where we start with the impurity modes being delocalized, which occurs at higher impurity concentration at $c_{A}=0.5$ and $V_{A}=0.9,\varepsilon=0.0,0.2,0.4$ In this case, as seen from Fig. 4-a) the impurity phonon modes are delocalized around $\omega=2$ to $4.5$ as depicted by a finite TDOS within this region. Our goal is to understand how $\varepsilon$ affects these delocalized phonons. As seen from Fig. 4-b) and Fig. 4-c), the delocalized phonons remain delocalized even when $\varepsilon$ is increased from 00 to $0.4$, although there are subtle changes in the mid-frequency region of the spectrum. We confirm this observation by examining the ratio of TDOS/ADOS, which remains almost unchanged as shown in Fig. 4-d).
@@
Finally, to further illustrate this effect, in Fig. 5, we examine a case with a highly delocalized spectrum which occurs at low disorder potential $V_{A}=0.5$. For such case, we observe that ADOS and TDOS are both finite across the entire frequency range for increasing values of $\varepsilon$, with the delocalized phonon spectrum being concentrated in the mid-frequency region for all values $\varepsilon$. Notably, while the TDOS shape shifts with increasing $\varepsilon$, the phonon modes remain delocalized. This is further confirmed by the near-constant ratio of TDOS/ADOS in Fig. 5-d), indicating minimal impact of $\varepsilon$ on phonon localization. Therefore, a low disorder potential favors highly delocalized phonon modes unaffected by $\varepsilon$ changes.
@@ IV Conclusions
In our study, we have introduced the multi-branch DCA and TMDCA formalisms to investigate the effects of disorder on the phonon spectrum in multi-branch systems. By comparing our results with exact ED results, we have found excellent agreement between the developed multi-branch DCA formalism and spectral functions for both binary and box disorder distributions. We have demonstrated the importance of the cluster extension of a single-site local CPA approximation to avoid averaging out essential non-local features in the spectrum. Although cluster DCA has advantages in incorporating non-local spatial correlations, it fails to capture Anderson localization due to the arithmetic averaging over disorder configurations. To address this limitation, we have developed the TMDCA multi-branch formalism, where a typical averaging ansatz replaces the arithmetic averaging in the self-consistency loop. Using the typical density of states as the order parameter for Anderson localization, we have investigated the impact of various vibrational multi-branch couplings ($\tau$ and $\varepsilon$) on phonon localization. Our study reveals that the multi-branch couplings in the model investigated have little effect on the Anderson transition of phonon localization in the presence of mass disorder. Specifically, our results demonstrate that localized states remain localized and extended states remain extended, largely regardless of the value of $\tau$ and $\varepsilon$. This does not rule out that multi-branch couplings in other models could be important for the localization of phonons. We leave, that question for future studies.
@@ acknowledgments
HT and WM acknowledge support by NSF DMR-1944974 grant. This work used Expanse at SDSC through allocation DMR 130036 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research also used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. A portion of this research (TB) was conducted at the Center for Nanophase Materials Sciences, which is a Department of Energy (DOE) Office of Science User Facility. In addition, we used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-05CH11231.

### v2

@@ I Introduction
Various numerical methods, including continuum field theory John et al. (1983), diagrammatic technique Kirkpatrick (1985), exact diagonalization(ED) Payton and Visscher (1967); Monthus and Garel (2010), replica methodJohn and Stephen (1983); John et al. (1983), multifractal analysis(MFA) Ludlam et al. (2003), transfer matrix method Pinski et al. (2012) have been employed to investigate phonon localization, particularly in simplified model systems. While real-space finite-size approaches provide unbiased benchmarks and controlled scaling, a practical constraint is the rapid growth in computational cost required to reach system sizes large enough to capture disorder effects accurately. For example, system sizes up to $18^{3}$ Monthus and Garel (2010), $48^{3}$  Ludlam et al. (2003), and $200^{3}$ Skipetrov and Beltukov (2018) are typically required, reflecting the exponential growth of the Hilbert space, the need for broad scaling windows and robust statistics, and the asymptotically large $L$ necessary to extract critical behavior via finite-size scaling.
As an alternative class to real space finite-size approaches, effective medium theories with controllable approximations provide a self-consistent framework for treating disorder. Within this framework, the coherent potential approximation (CPA) Soven (1967); Elliott et al. (1974) serves as the canonical single site scheme, where a translationally invariant medium is determined self consistently from disorder averaged quantities. Building on CPA, cluster extensions including the dynamical cluster approximation (DCA) Jarrell and Krishnamurthy (2001); Ekuma et al. (2014); Terletska et al. (2014); Ekuma et al. (2015); Zhang et al. (2015, 2016) embed a finite cluster of size $N_{c}$ in a self-consistently determined effective medium, which incorporates short-range nonlocal correlations and improves accuracy for both electronic and vibrational models. For lattice dynamics, the itinerant coherent potential approximation (ICPA) Ghosh et al. (2002) extends CPA to include mass and force constant disorder and has been used to describe vibrational spectra in realistic binary alloys.
Although CPA and DCA accurately reproduce disorder-averaged spectra, they do not provide a description of Anderson localization Ekuma et al. (2014). The typical-medium dynamical cluster approximation (TMDCA) addresses this by incorporating geometric averaging of the local density of states within the cluster self-consistency, so that the typical density of states (TDOS) serves as an order parameter that vanishes at localization Ekuma et al. (2014). In TMDCA, a finite cluster of size $N_{c}$ is embedded in a self-consistent effective medium; the approximation is controllable since $1/N_{c}$ acts as a small parameter, and increasing $N_{c}$ systematically adds nonlocal correlations, yielding convergence toward the exact result. This construction accesses thermodynamic-limit disorder averages without explicit finite-size scaling or prohibitively large real-space lattices. For the three-dimensional Anderson tight-binding model, TMDCA reproduces the critical disorder strength $W_{c}$ and the localization phase diagram for electron localization with box disorder in excellent agreement with transfer-matrix method (TMM) results using moderate clusters (e.g., $N_{c}=92$Ekuma et al. (2014). In practice, the Green’s function-based TMDCA offers computational efficiency and flexibility that make it particularly well-suited for investigating complex disordered systems. These include systems with multiple orbitals, long-range disorder potentials, electron-electron interactions, and features often inaccessible to conventional finite-size approaches Terletska et al. (2018); Zhang et al. (2016); Ekuma et al. (2014); Terletska et al. (2014); Ekuma et al. (2015); Sen et al. (2016); Tam et al. (2021a); Terletska et al. (2021); Tam et al. (2021b).
While Anderson localization of electrons arises from interference of scattered electronic wave functions, the corresponding phenomenon for phonons occurs in the vibrational spectrum of disordered lattices. In such systems, disorder effects are most pronounced at high frequencies (short wavelengths), whereas low-frequency acoustic modes remain extended even under strong disorder. Numerical studies using finite-size methods, such as ED Monthus and Garel (2010), MFA Ludlam et al. (2003), and finite-size scaling  Skipetrov and Beltukov (2018) on large system sizes, together with experimental observations Howie et al. (2014), uch delocalization of long-wavelength phonons. These studies further demonstrate that the localization transition of high-frequency modes belongs to the same universality class as electronic Anderson localization Monthus and Garel (2010); Ludlam et al. (2003); Skipetrov and Beltukov (2018). Consequently, the effective medium TMDCA, which is well established for electronic systems, can be extended to study phononic systems Mondal et al. (2017). In contrast to the very large lattices required for the finite-size studies, TMDCA reproduces the phonon localization phase diagram in quantitative agreement with transfer-matrix-method (TMM) results using modest clusters (e.g., $N_{c}=4^{3}$) while retaining computational efficiency Mondal et al. (2017). TMDCA has been established to be a successful method for studying Anderson localization of phonons in all disorder limits Mondal et al. (2017). However, previous TMDCA implementations for phononic systems with mass Mondal et al. (2017) and force-constant Mondal et al. (2019) disorder have been restricted to single-branch models, i.e., vibrations in a single polarization direction.
@@
Despite this progress, modeling phonon localization in realistic materials remains challenging. An effective numerical method must construct material-specific force-constant models and solve the resulting lattice-dynamical problem. The absence of lattice translational symmetry in disordered solids and the breakdown of the Bloch theorem make this task difficult to solve straightforwardly. Moreover, force-constants calculations are computationally expensive compared to electronic case hopping calculations, and maintaining the force-constants sum rule can be challenging, especially in the presence of disorder. The inclusion of vibration directions adds another layer of complexity to the problem due to the correlations, where the vibration of an atom at site $R_{l}$ in the $\alpha$ direction is coupled with the vibration of an atom at site $R_{l^{\prime}}$ in the $\beta$ direction. Furthermore, accurately capturing Anderson localization in such systems is challenging, as it often requires the calculation of two-particle transport quantities or finite-size scaling analysis that can be computationally expensive, particularly for complex realistic models.  Zhang et al. (2017). Nonetheless, first-principles-based workflows have recently begun to tackle phonon localization in specific materials systems. Magdău and Ackland (2017); Zhang et al. (2022)
@@
In this work we extend TMDCA to treat phonons with three polarization directions, enabling more realistic modeling of vibrational spectra and localization within an effective-medium framework. This development builds on earlier TMDCA studies of phonon localization in single-branch models with mass disorder Mondal et al. (2017) and with force-constant disorder Mondal et al. (2019). Here we formulate a multi-branch ( with three directions of vibrations) TMDCA for mass-disordered lattices. We validate the developed formalism against exact-diagonalization benchmarks and appropriate limiting cases. Applying our developed multi-branch formalism, we investigate how the vector nature of vibrations affects the spectra and localization of phonons. After comparing the results of single-branch and multi-branch situations, we found that the Anderson transition of phonons is not significantly influenced by the vector nature of the phonons for the specific model considered in this work. Overall, the present study provides further understanding of the localization of vector phonons and brings the TMDCA one step closer to being an efficient numerical tool for studying Anderson localization of phonons in materials from first principles.
@@ B. Green’s function formalism
$\langle\langle$$\rangle\rangle$ stands for the expectation value and time ordering between two displacement ($u^{\alpha}_{i}(l)$) and ($u^{\beta}_{j}(l)$) operators.
@@ 4. We then construct the cluster Green’s function for a given disorder configuration $V$:
here $\mathcal{D}$, D{D} are $N_{b}N_{c}\times N_{b}N_{c}$ matrices in the real space of the cluster. In our analysis, we consider a binary alloy disorder described by the following distribution function
@@
where <><...> indicates an averaging over disorder configurations.
@@ D. Typical medium dynamical cluster approximation for phonons
$D_{\scriptscriptstyle{\rm typ}}^{\alpha\beta}(\mathbf{K},\omega)={\mathcal{P}}\int d\omega^{\prime}\frac{{\rho^{\alpha\beta}_{\rm typ}({\mathbf{K}},\omega^{\prime})}}{\omega^{2}-\omega^{\prime 2}}-i\frac{\pi}{2\omega}{{\rho^{\alpha\beta}_{\rm typ}(\mathbf{K},\omega}})\,$, (22)
@@ 2.1 Multi-branch force constants model
$\Phi^{\alpha\beta}(\vec{R}_{l},\vec{R}_{l})$ $6\times\tau$
@@ A. Benchmarking of the DCA and ED methods
We start the discussion of our results by first comparing our developed multi-branch cluster DCA method with the ED results for a binary isotopic alloy system in three dimensions. This benchmarking analysis aims to evaluate the accuracy and effectiveness of our developed multi-branch cluster DCA approach in capturing the system’s behavior. For this, in Fig. 1, we compare the average density of states (ADOS) calculated for an impurity concentration of $c_{A}=0.2$ and a disorder potential of $V=0.9$, for a fixed value of $\tau=0.3$ and varying values of $\varepsilon$. First, on panel Fig. 1-a), we consider the case of $\varepsilon=0.0$, where we compare the DCA and ED results, and observe a perfect agreement between two methods. For the finite values of $\tau$, we anticipate the emergence of two impurity modes in the higher frequency regions. Specifically, we expect a doubly degenerate impurity band at the lower end of the high-frequency region and a single impurity band at the higher end, corresponding to the eigenvalues of the $3\times 3$ onsite matrix $\Phi^{\alpha\beta}(l,l)$: $1-\tau,1-\tau,1+2\tau$ as explained in section  II. As shown in Fig. 1, the two impurity modes around frequency regions $2<\omega<4$ and $4<\omega<6$ are well captured by our DCA method and are in perfect agreement with the ED results.
@@ C. $\varepsilon\neq 0$ and $\tau\neq 0$: inter-branch coupling case
In Fig. 3, to demonstrate the impact of inter-branch coupling on phonon localization, we plot the ADOS and TDOS for increasing values of $\varepsilon=0.0,0.2,0.4$ at a fixed disorder potential $V_{A}=0.9$ and small impurity concentration of $c_{A}=0.2$. Beginning with the branch-decoupled case ($\varepsilon=0.0$), as shown in Fig. 3-a), we observe strong localization for high-frequency phonon modes within the frequency range of $\omega=2.5$ to $\omega=6.0$. This is evident from the TDOS approaching zero while the ADOS remains finite. The results for finite $\varepsilon=0.2$ and $\varepsilon=0.4$ are shown in Fig. 3-b) and Fig. 3-c). The data indicate a negligible effect of $\varepsilon$ on the localized phonons, although subtle changes are observed in the host modes. Specifically, the host mode becomes slightly broader and more delocalized with increasing $\varepsilon$. However, high-frequency phonons remain localized even as $\varepsilon$ increases. This is clearly illustrated in Fig. 3-d), where the ratio of TDOS/ADOS remains nearly constant across the range of considered $\varepsilon$ values (0 to $0.4$).
@@
Next, we consider an opposite scenario where we start with the impurity modes being delocalized, which occurs at higher impurity concentration at $c_{A}=0.5$ and $V_{A}=0.9,\varepsilon=0.0,0.2,0.4$ . In this case, as seen from Fig. 4-a) the impurity phonon modes are delocalized around $\omega=2$ to $4.5$ as depicted by a finite TDOS within this region. Our goal is to understand how $\varepsilon$ affects these delocalized phonons. As seen from Fig. 4-b) and Fig. 4-c), the delocalized phonons remain delocalized even when $\varepsilon$ is increased from 0 to $0.4$, although there are subtle changes in the mid-frequency region of the spectrum. We confirm this observation by examining the ratio of TDOS/ADOS, which remains almost unchanged as shown in Fig. 4-d).
@@
To further illustrate this effect, in Fig. 5, we examine a case with a highly delocalized spectrum which occurs at low disorder potential $V_{A}=0.5$. For such case, we observe that ADOS and TDOS are both finite across the entire frequency range for increasing values of $\varepsilon$, with the delocalized phonon spectrum being concentrated in the mid-frequency region for all values $\varepsilon$. Notably, while the TDOS shape shifts with increasing $\varepsilon$, the phonon modes remain delocalized. This is further confirmed by the near-constant ratio of TDOS/ADOS in Fig. 5-d), indicating minimal impact of $\varepsilon$ on phonon localization. Therefore, a low disorder potential favors highly delocalized phonon modes unaffected by $\varepsilon$ changes.
D. Mobility edge analysis for box disorder
[…]
Figure 7: The mobility edge trajectories of the phonon spectrum in three dimensions for a box distribution.
Finally, we investigate the influence of coupling parameters on the mobility edge Anderson (1958); Abrahams et al. (1979); Lagendijk et al. (2009); Chang et al. (2025), a critical quantity that separates localized and delocalized states in three-dimensional disordered systems. To this end, we consider a box disorder distribution defined as $P_{V}(V_{l})=\Theta(V-|V_{l}|)/2V$, where $V_{l}=(1-M_{l}/M_{0})$, and $0\leq V\leq 1$, where $l$ is the lattice site.
We identify the mobility edge as the high-frequency band edge of TDOS($\omega$), which marks the transition from extended to localized vibrational states. Fig. 6 presents the TDOS for both the decoupled case ($\tau=0,\varepsilon=0$) and a finite coupling case ($\tau=0.2,\varepsilon=0.4$) for various strengths of the disorder $V$. As seen in the figure, the TDOS is increasingly suppressed with higher disorder $V$ for both coupling regimes. Notably, in the presence of finite branch coupling, the TDOS broadens, and the high-frequency band edge shifts to higher energies, indicating a modification of the mobility edge.
[…]
We now turn our discussion to the coupled scenario with $\tau=0.2,\varepsilon=0.4$. As shown in Fig. 7, finite inter-branch coupling shifts the high-frequency edge of the TDOS toward higher values for any given disorder strength $V$. For example, at $V=0.4$, the right band edge of the TDOS shifts from $\omega^{2}=3.33$ to $\omega^{2}=4.70$, and at $V=0.9$, it shifts from $\omega^{2}=2.40$ to $\omega^{2}=3.61$. The overall shape and character of the mobility edge trajectory still remain largely unchanged in the finite inter-branch coupling case. In particular, the re-entrant behavior of the mobility edge around $V=0.5$ persists at these coupling parameters. Overall, our results suggest that coupling between different vibrational branches does not significantly alter the very nature of the Anderson transition in phonons with mass disorder for the box disorder with re-entrance behavior being present in both coupled and decoupled cases.
@@ IV Conclusions
In our study, we have introduced the multi-branch DCA and TMDCA formalisms to investigate the effects of disorder on the phonon spectrum in multi-branch systems. By comparing our results with exact ED results, we have found excellent agreement between the developed multi-branch DCA formalism and spectral functions for both binary and box disorder distributions. We have demonstrated the importance of the cluster extension of a single-site local CPA approximation to avoid averaging out essential non-local features in the spectrum. Although cluster DCA has advantages in incorporating non-local spatial correlations, it fails to capture Anderson localization due to the arithmetic averaging over disorder configurations. To address this limitation, we have developed the TMDCA multi-branch formalism, where a typical averaging ansatz replaces the arithmetic averaging in the self-consistency loop. Using the typical density of states as the order parameter for Anderson localization, we have investigated the impact of various vibrational multi-branch couplings ($\tau$ and $\varepsilon$) on phonon localization. Our study reveals that the multi-branch couplings in the model investigated have little effect on the Anderson transition of phonon localization in the presence of mass disorder. Specifically, our results demonstrate that localized states remain localized and extended states remain extended, largely regardless of the value of $\tau$ and $\varepsilon$. We also show that, for the box disorder distribution, inter-branch coupling does not alter the re-entrant behavior of the mobility edge, preserving the overall nature of the Anderson transition observed in the decoupled case. These results do not rule out that multi-branch couplings in other models could be important for the localization of phonons. We leave, that question for future studies.
@@ acknowledgments
WM acknowledge support by NSF DMR 1944974 grant (method development). HT and TB acknowledge support by the U.S. Department of Energy, Office of Science, under award number DE-SC0025748 grant. This work used Expanse at SDSC through allocation DMR 130036 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research also used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. A portion of this research (TB) was conducted at the Center for Nanophase Materials Sciences, which is a Department of Energy (DOE) Office of Science User Facility. In addition, we used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-05CH11231.